a car rental for $85 per day plus $0.35 for every mile driven more than 200 miles. If the car was driven 300 miles and the rental paid was $205, for how many days was the car rented?
Answer by ankor@dixie-net.com(22740)

a car rental for $85 per day plus $0.35 for every mile driven more than 200 miles.
If the car was driven 300 miles and the rental paid was $205, for how many days was the car rented?
:
let d = the no. of days the car was rented
85d + .35(300 - 200) = 205
85d + .35(100) = 205
85d + 35 = 205
85d = 205 - 35
85d = 170
d = 170/85
d =2 days